Find the extreme values of 'f' on the region described by the inequality.

22. f (x, y) = 2x2 + 3y2 – 40 – 5, x2 + y2 < 16
Answer #1

tu (y) = 2x² + 3y 40.5 g (2,4)= 16- 4² ye fu (ny) Un-4 fy(x, y)- by Ju (x, y) = - 2x gy (x, y) = -24 fx= dgn le tye day 74n-y